Journalist Report – May 14th

Hello Mission Control,

First off, thank you Ben for saving us from bad hygiene with a kind soap refill–we really appreciate it :D! Today we had a day with our main theme being teamwork. We were awake early and Avery, Rishabh and Kristina were ready to suit up the three of us up for our EVA. The materials trio drove out around 8:30AM and had a great time! We were able to test out the rovers that mission control requested as well as get some sample (which was not on the plan but we carried supplies in case we found something interesting!). We then got back and I dropped off the sample to the Science Dome. Me, Aravind, and Noah then helped the other three suit up for their EVA. I stayed on comms and they had a successful round as well! We later had a meeting to plan out two more EVAs and created a schedule for the day Spencer Joseph from FOX is here. We are planning in advance to time the content with the research we will be doing. The astronomers are hard at work right now analyzing some content from the recent solar flares tonight since the weather was unfortunately not permitting observing and we will have dinner soon.

Bevonauts out!

Prakruti "Pari" Raghunarayan (Crew Commander & Journalist)

Sol Summary – May 14th

Crew 299 Sol Summary 05-13-2024
Sol: 2
Summary Title: The First Adventure
Author’s Name: Prakruti "Pari" Raghunarayan
Mission Status: Active
Sol Activity Summary:

Pari, Aravind, Noah:
Today, Noah did a morning check on the radishes. Then we went on our first test EVA to Marble Ritual. We tested comms and took Perseverance as requested by Mission Control. We are pleased to report that the rover is working fine. We took Curiosity as well. When we got to Marble Ritual we parked and walked up a hill to test out our mobility in the suits. We were able to hike, take samples of some rocks we found that we believe will be interesting to look at the composition as the map indicated some sort of river bed that we would further like to investigate. We were able to drive back and end the mission smoothly. I (Pari) was also on comms when Kristina, Avery, and Rishabh went for their test EVA as well.

Avery, Kristina, Rishabh:
For Sol 2, we went on our test EVAs to Marble Ritual, taking photos for geological reference and social media outreach. Then, we had tacos for lunch and cleaned up the Hab. The weather was not suitable for observing, so Kristina and Avery worked on processing yesterday’s images. Later, Pari gave a debrief outlining a plan of action for the rest of the day, tomorrow, and the media visit

Sol Summary – May 13th

Crew 299 Sol Summary 05-13-2024
Sol: 1
Summary Title: Pilot
Author’s Name: Prakruti Raghunaryan
Mission Status Active:
Sol Activity Summary:

Kristina and Avery Summary:

First thing, we had a morning debrief to discuss a plan of action. Then, we received the laptop and did a secondary round of laptop repair attempts. We found the battery was dead, and the boot drive wasn’t connected. After, we fitted flight suits with the whole crew and chose which ones were best suited for each of us for our future EVAs. Once that was completed we broke for lunch and planned tomorrow’s EVAs. Command got back to us on the issues with our laptop. For the third time, we troubleshooted astronomy laptop logistics and got it working by plugging the boot drive into Avery’s laptop. With the current solar storms, we will try to get the Musk Observations working before sunset. One possible issue is that since the astronomy laptop could only connect to one USB, it may not connect to the telescope, as one USB port is currently being used to connect to Avery’s laptop. Then, we did our psychological evaluations and dinner.

Pari, Aravind, and Noah:

We started the day with a morning debrief to mark the start of our simulation. We created a plan for the day with the rest of our crew mates and headed to the science dome. We exfoliated four different types of samples: unsorted Martian soil, fine Martian soil, potting soil, and vermicompost. We also checked on the radishes and watered them. Then we went back and suited up in our flight suits. We did the adjustments for the EVA suits (for the two piece and one piece). We broke for lunch and took a short break after. Next, we had a meeting to plan the EVAs for tomorrow and the day after. We planned two EVAs for tomorrow and are likely getting more external sample to test out from the river bank. Then we went back into the science dome and exfoliated more of the unsorted and fine soil onto the PDMS. We also tested out the microscope. We created a LinkedIn page to showcase our current progress and are waiting for our crew members to finish up the astronomy project. Now we will break for dinner and do the psychological evaluations.

Crew Engineer did maintenance and checked the suits before try ons.

Journalist Report – May 13th

Hello Mission Control,

Today we had a relatively successful first day. Though it involved some problems, we overcame them and found solutions. I started the day by holding a team debrief where we went over today’s tasks and crew goals. I joined the crew engineer, Rishabh, to do maintenance. We replaced a few zip ties, measured the water, and checked voltage on the suits. Then, we split into our subgroups for the day. Me, Noah (Crew Biologist), Aravind (Crew Chemist) worked on the science dome samples. Kristina (HSO), Avery (XO/Astronomer), and Rishabh worked on rebooting the astro laptop. We later met to get our suits adjusted and tested for our EVAs tomorrow! We got some pictures and are very excited! Additionally, the three of us went back to the science dome to continue after lunch and the astronomers continued their work and persisted despite issues faced closing the dome. Overall, we worked great as a team today and it was enjoyable!

See you tomorrow 🙂

Pari Raghunarayan (Crew Commander & Journalist)

Operations Report – May 13th

Crew 299 Operations Report 05-13-2024

SOL: 1

Name of person filing report: Rishabh Pandey (Engineer)

Non-Nominal Systems: Suit 9 (broken visor), Suit 11 (fan issue), Observatory Lower Deck (broken)

Notes on non-nominal systems: Issues with both suits were noted before our mission, both will not be used during our mission. Issue with the Observatory Lower Deck was noted before our mission, will be manually opened and closed.

No Rovers used.
General notes on rovers: NA

Summary of Hab operations:
WATER USE: NA, first day of measurements
Water (static tank): 530 (estimated) first day of measurement
Static tank pipe heater: off
Static tank heater: off
Toilet tank emptied: yes

Summary of Internet usage: Catch up on Earthly news, email responses, code learning + development

Summary of suits and radios:
Radios 1,2,3,4,5 and 6 were used
EVA Suits 7,8, and 10 were used (worn for suit adjustments and taken off)

Summary of GreenHab operations: NA, GreenHab out of operation

Summary of ScienceDome operations: Soil exfoliation and checking on radishes

Summary of RAM operations:
The astronomy laptop had a failed boot drive, the drive was extracted and successfully booted on another Windows machine. All software from the Astronomy laptop is usable on a new interface.
Fixed missing zip ties: 4 replaced on the tunnel to the RAM, 4 replaced on the main tunnel, 3 replaced on the tunnel to ScienceDome, 3 replaced on the tunnel to Observatory
Tools Used: Multimeter (suit power check), Measuring tape: Initial water level check, Scissors to cut zip ties

Summary of observatory issues: Temporary issue with the upper deck failing to close as a result of a broken lower deck. Needed to manually press down the lower shutter limit switch in order to trigger the closing motor and mechanism for the upper deck.

Supplemental Operations Report – May 19th

Date: 5/19/2024
Name of person filing report: Ben Stanley
Reason for Report: Routine
Non-Nominal Systems: Hab wall heater, solar observatory, perseverance parking brake & battery. EVA suit 9, Solar Power inverters slave 1&2, astronomy laptop dead

Power system:
Solar: Malfunctioning inverters still out for repairs.
SOC Last 24 hours: Max 100%; Min 90%; Avg 97.3%.
VDC Last 24 hours: Max 58.27V; Min 49.17V; Avg 53.86V.
Generator run time: 6139.7. Technician diagnosed problem as bad starter. They didn’t have a replacement so one was ordered. Ben installed on 5/16 and main generator is working again.

Propane Readings:
Station Tank: 77%
Director Tank: 71%
Intern Tank: 79%
Generator Tank: 74%

Hab Static Tank – 359 gallons
GreenHab – 0 gallons
Outpost tank – 370 gallons
Science Dome – 0 gallons
Septic Tanks emptied: No

Sojourner rover used: No
Hours: 200.4
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: No
Notes on Rovers: Perseverance is now charging only up to about 60%. Charging error appears from time to time. Rovers moved back to area in front of RAM now that the main generator is back online.

ATV: ATV’s Used: None. Nothing to report.

Hab Car used and why, where: To Hanksville for supplies.
Crew Car used and why, where: No.
General notes and comments: Ignition problem persists in habcar. Can get it to start by resetting into park and trying to pull the shifter as high as possible.

Summary of Internet: All nominal.

EVA suits and radios:
Suits: Suit 9 has damaged visor and put aside until end of the season.
Comms: Nominal.

Campus wide inspection, if action taken, what and why: Solar Power still at 4.4kW. Main generator repaired and being used to charge rovers as needed.

Summary of Hab Operations: Wall heater is not functioning. Diagnostic is pending.
Summary of GreenHab Operations: Shut down for summer.
Summary of SciDome Operations: Inverters still out for repairs.

Summary of Observatories Operations: Lower shutter is decoupled and fastened by bungee cord. Astronomy laptop is dead.

Summary of RAM Operations: All nominal
Summary of Outpost Operations: All nominal
Summary of Health and Safety Issues: One crew member had what appears to have been a cold. They have recovered and all is nominal.

EVA Report – May 14th

Crew 299 EVA 1 Report 5-14-2024

EVA #1

Author: Rishabh Pandey

Purpose of EVA: Test EVA up to Marble Ritual to get certified.

Start Time: 8:30AM

End Time: 9:25AM

Narrative: The EVA team took curiosity and perseverance rovers up to Marble Ritual for their first-ever EVA. The crew took some fun pictures to be uploaded to social media and for the journalist. After walking around the area and climbing the hill for a short period, the crew returned to the HAB.

Destination: Marble Ritual

4251000 5187500 (Marble Ritual)

EVA Participants: Prakrut (Crew Commander & Journalist), Noah (Biologist), Aravind (Chemist)

Road(s) and routes per MDRS Map: Cow Dung Road

Mode of travel: Driving Opportunity + Walking

Copyright © The Mars Society. All rights reserved. | Main Site